The book is Algebra, 3/e, by K. Elayn Martin-Gay!

Determine whether each of the following is an equation or an expression. If is an Equation, solve it for its variable. If its an expression, perform the indicated operation.
1. 1 2
- + -
x 3
I know the we need to find a LCD but i not sure wich one would be i got x(3)for my LCD????
and for my answer i got 3+2x over 3x
...but i not sure if thats correct...Please Help...if you can help me please explain me why, and how can i get it..so i can understand...Thanks

Answer by checkley75(3666)

1.1/X+2/3=(3*1.1+2X)/3X=(3.3+2X)/3X